# Quantum Machine Learning: from NISQ to Fault Tolerance

### v1

@@ II.2 Quantum neural tangent kernel
$d\bar{}\hskip 1.00006ptz_{i;\delta}=d\bar{}\hskip 1.00006pt\varepsilon_{i;\delta}\equiv\frac{\partial z_{i;\delta}}{\partial\theta_{\ell}}\left(\mathcal{D}\theta_{\ell}\right)=\frac{dz_{i;\delta}}{d\theta_{\ell}}\left(-\eta\frac{\partial\mathcal{L}_{\mathcal{A}}}{\partial\theta_{\ell}}\right)=-\eta\sum_{j,\alpha,\ell}\varepsilon_{i^{\prime};\alpha}\frac{\partial\varepsilon_{i;\delta}}{\partial\theta_{\ell}}\frac{\partial\varepsilon_{i^{\prime};\alpha}}{\partial\theta_{\ell}}=-\eta\sum_{j,\alpha}\varepsilon_{i^{\prime};\alpha}K^{i,i^{\prime}}_{\delta,\alpha}~{},$ (5)
The variational angles $\theta_{\ell}$ are then updated step by step, labeled by $t$, according to the derivative of the loss function with respect to $\theta_{\ell}$. $\eta$ is defined as the learning rate, which is assumed to be a small (less than 1), positive number. The learning rate determines the size of the step that we take to update our parameters. One can calculate the update of the loss function $\mathcal{L}$ and prove that it decreases indefinitely every time we update our parameters, which is also a fact in classical machine learning Roberts et al. (2022). In Eqn. (5) it is easy to show that the update of the output is mathematically in the exact same form as the update of the residual training error $d\bar{}\hskip 1.00006pt\varepsilon_{i;\delta}$. One then defines the quantum neural tangent kernel (QNTK):
$K^{i,i^{\prime}}_{\delta,\alpha}\equiv\sum_{\ell}\frac{\partial\varepsilon_{i;\delta}}{\partial\theta_{\ell}}\frac{\partial\varepsilon_{i^{\prime};\alpha}}{\partial\theta_{\ell}}~{}.$ (6)
@@
However, the final assumption is not always true. To wit, with the expression of the output $z_{i,\alpha},$ one can analytically calculate the QNTK and prove that the QNTK is highly fluctuating in the variational angles $\bm{\theta}$, rather than being a constant Liu et al. (2022c). This fluctuation implies that we have obtained representation learning. In the terminology of the kernel method (NTK) of classical machine learning, it means that the feature map itself is changing. This indicates that we are extracting features from all of our training data to update our output, and we cannot obtain a closed-form solution for $\varepsilon(t)$ in this case.
@@
Figure 3: Demonstration of the evolution of the loss function in the continuous limit Liu et al. (2022c). The right diagram depicts the closed-form solution when the QNTK remains constant. In contrast, the left diagram illustrates a scenario where the QNTK is nonlinear but nearly constant. In this case, representation learning is achieved, and the gradient descent equation can be addressed perturbatively.
An instance where the QNTK remains nearly constant is referred to as “lazy training”. This occurs when the variational angles are nearly unchanged Liu et al. (2022c); Chizat et al. (2020): $\theta_{\ell}=\theta_{\ell}^{*}+\delta\varphi~{}.$ If $\theta_{\ell}$ remains approximately constant, indicating lazy training with minimal movement, the QNTK tends to be nearly constant as well. In such cases, the QNTK expression can be expanded with respect to $\theta_{\ell}$ for predictable dynamics, akin to classical scenarios Roberts et al. (2022).
@@
$\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu=\frac{L\operatorname{Tr}(\mathcal{O}_{i}^{2})}{N^{2}}\propto L~{}.$ (7)
@@
Nonetheless, this model remains useful, as even when not operating at the frozen limit, it provides an approximate means to gauge the speed of gradient descent. For example, to determine which variational ansatz runs the fastest without actually executing them, one can calculate their QNTK and select the one with the largest value. It also enables the estimation of operation time and the number of iterations required for the gradient descent program using Eqn. (7), facilitating hardware design improvements. Moreover, some late-time behaviors of QNTK also been recently captured in Zhang et al. (2023).
II.3 Quantum landscapes and barren plateaus
@@
The process of training parameters in the Quantum Machine Learning (QML) model often involves minimizing a loss function and traversing a non-convex landscape in search of its global minimum, as extensively discussed in Cerezo et al. (2022a). Quantum landscape theory, as outlined in Arrasmith et al. (2022), seeks to comprehend the properties of QML landscapes and explores strategies for engineering them. Notably, local minima and barren plateaus are focal points of investigation within quantum landscape theory
@@
Akin to classical ML Cerezo et al. (2022a), the quantum loss landscape may contain numerous local minima. In a manner similar to classical cases, the overall non-convex optimization can become NP-hard Bittel and Kliesch (2021). Various approaches have been proposed to mitigate issues related to local minima. For instance, variable structure Quantum Neural Networks (QNNs), as introduced in Bilkis et al. (2023); LaRose et al. (2019), dynamically adjust the model’s prior by expanding and contracting during optimization, transforming certain local minima into saddle points. Additionally, there is evidence of the overparametrization phenomenon in QML Kiani et al. (2020); Larocca et al. (2023). In this scenario, a computational phase transition occurs as the optimization progresses, leading to the disappearance of spurious local minima when the number of parameters surpasses a critical value.
@@
Not only are local minima a concern in QML, but there is also the intriguing phenomenon of a barren plateau in quantum landscapes Cerezo et al. (2022a, 2021); Arrasmith et al. (2021); Holmes et al. (2022). A barren plateau is characterized by the loss landscape becoming, on average, exponentially flat concerning the problem size. In this scenario, the valley housing the global minimum exponentially diminishes with the problem size, forming what is known as a narrow gorge Arrasmith et al. (2022). Consequently, exponential resources, such as an increased number of shots, are needed to traverse through the landscape. This impact on resource requirements complicates the QML algorithm’s complexity and can potentially undermine quantum speedup, as quantum algorithms typically aim to avoid the exponential complexity associated with classical algorithms.
@@
The barren plateau phenomenon was first studied in deep hardware-efficient QNNs McClean et al. (2018), where they arise due to the high expressivity of the model Arrasmith et al. (2021); Cerezo et al. (2022a). By making no assumptions about the underlying data, deep hardware-efficient architectures aim to solve a problem by being able to prepare a wide range of unitary evolutions. In other words, the prior over the hypothesis space is relatively uninformed. Barren plateaus in this unsharp prior are caused by ignorance or the lack of sufficient inductive bias, and therefore a means to avoid them is to input knowledge into the construction of the QNN - making the design of QNNs with good inductive biases for the problem at hand a key solution. Various strategies have been developed to address these barren plateaus, such as clever initialization Verdon et al. (2019), pretraining, and parameter correlation Holmes et al. (2022). These are all examples of adding a sharper prior to one’s search over the over-expressive parameterizations of hardware-efficient QNNs.
@@
Other mechanisms have been linked to barren plateaus Cerezo et al. (2022a). Simply defining a loss function based on a global observable (i.e., observables measuring all qubits) leads to barren plateaus even for shallow circuits with sharp priors McClean et al. (2018), while local observables (those comparing quantum states at the single-qubit level) avoid this issue McClean et al. (2018); Uvarov and Biamonte (2021). The latter is not due to bad inductive biases but rather to the fact that comparing objects in exponentially large Hilbert spaces requires an exponential precision, as their overlap is usually exponentially small.
@@
While entanglement is one of the most important quantum resources for information processing tasks in quantum computers, it can also be detrimental for QML models Cerezo et al. (2022a). QNNs (or embedding schemes) that generate too much entanglement also lead to barren plateaus Sharma et al. (2022); Marrero et al. (2021); Patti et al. (2021). Here, the issue arises when one entangles the visible qubits of the QNN (those that one measures at the QNN’s output) with a large number of qubits in the hidden layers. Due to entanglement, the information of the state is stored in non-local correlations across all qubits, and hence the reduced state of the visible qubits concentrates around the maximally mixed state. This type of barren plateau can be solved by taming the entanglement generated across the QNN.
II.3.1 Influence of background noise
@@
The existence of hardware noise during quantum computations is a distinctive feature of NISQ computing Cerezo et al. (2022a). Despite this reality, many QML studies overlook noise in analytical calculations and numerical simulations, while still asserting near-term compatibility of the methods. It is imperative to consider the impact of hardware noise in QML analyses, especially for those aiming to achieve quantum advantage with currently available hardware Cerezo et al. (2022a).
@@
Noise introduces errors in the information as it propagates through a quantum circuit, particularly affecting deeper circuits with longer run-times Cerezo et al. (2022a). This impact extends to all aspects of quantum models, including the dataset preparation scheme and circuits used for computing quantum kernels. In the case of Quantum Neural Networks (QNNs), noise can impede their trainability Wang et al. (2021a, b), leading to noise-induced barren plateaus where relevant features are exponentially suppressed with increasing circuit depth. The consequence is a deformation of the model’s inductive bias and an effective reduction in the dimension of the quantum feature space. Despite the significant influence of quantum noise, its effects remain largely unexplored, especially regarding its impact on the classical simulability of QML models Deshpande et al. (2022); Hakkaku et al. (2022); Cerezo et al. (2022a).
@@
Addressing issues induced by noise may necessitate Cerezo et al. (2022a): (1) reducing hardware error rates, (2) implementing partial quantum error correction Bultrini et al. (2023), or (3) utilizing QNNs that are relatively shallow (i.e., with depth growing sub-linearly in the problem size) Wang et al. (2021a), such as Quantum Convolutional Neural Networks (QCNNs). While error mitigation techniques Temme et al. (2017); Czarnik et al. (2021); Endo et al. (2021) can enhance the performance of QML models in the presence of noise, they may not completely resolve noise-induced trainability issues Wang et al. (2021b). Another approach involves engineering QML models with noise-resilient properties LaRose and Coyle (2020); Cincio et al. (2021); Sharma et al. (2020), such as ensuring that the position of minima remains unchanged despite noise.
@@
However, another study suggests that noise in VQAs might be beneficial for their performance. Drawing inspiration from classical machine learning, where noise is intentionally introduced to enhance the performance of the gradient descent algorithm, this phenomenon is akin to stochastic gradient descent. It has been observed that gradient descent tends to perform better under the influence of noise, as the noise helps in avoiding saddle points that could hinder the algorithm’s progress Jain and Kar (2017); Jin et al. (2019). In the absence of noise, the algorithm may converge directly to the saddle point. However, when noise is introduced, perturbations around the points prevent convergence towards these unstable fixed points.
@@
In the study Liu et al. (2022a), this discussion was extended to quantum noise. Through experiments, the researchers found that the noisy case outperformed the noiseless counterparts, providing a toy model illustration to support their argument. This suggests that, while implementing quantum error correction is crucial for making quantum devices fault-tolerant, in certain QML applications, eliminating all noise may not be necessary. As long as the noise remains below a certain threshold, it may prove not detrimental but rather helpful for QML programs.
II.3.2 Laziness
@@
$\theta_{\ell}(t+1)-\theta_{\ell}(t)=-\eta\frac{\partial\mathcal{L}}{\partial\theta_{\ell}}\equiv\delta\theta_{\ell}~{}.$ (9)
@@
The observation McClean et al. (2018) is that if our variational ansatz is highly random, according to the $k$-design integral formula Roberts and Yoshida (2017); Liu (2020, 2018); Cotler et al. (2017), the derivative of the loss function is generally suppressed by the dimension of the Hilbert space $N$. This might lead to a situation where the variation of the loss function during gradient descent is very small, denoted as $\delta\mathcal{L}\equiv\mathcal{L}(t+1)-\mathcal{L}(t)\ll 1$ for the step $t$. An alternative theoretical term used to describe the quantum barren plateau, based on the large suppression in Eqn. (9), is laziness Liu et al. (2022b). In the quantum context, the suppression comes from the dimension of the Hilbert space, while in the classical case, the suppression is associated with the width of classical neural networks.
@@
To be more precise, laziness refers to small $\delta\theta_{\mu}$, and the barren plateau refers to small $\delta\mathcal{L}$. The authors of Liu et al. (2022b) demonstrate that laziness may not necessarily imply the quantum barren plateau. This insight is derived from the perspectives of overparametrization theory and representation learning theory through the QNTK Liu et al. (2022c). In the context of quantum neural networks, overparametrization refers to the condition $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu\approx\mathcal{O}(1)$ in Eqn.(7).
@@
The paper Liu et al. (2022b) emphasizes that in variational circuits with a sufficiently large number of trainable angles, gradient descent dynamics can still be efficiently executed, even in the presence of the exponential suppression of variational angle updates (laziness). On the other hand, the only limitation is the precision, where noises might affect the performance significantly. The authors also highlight that laziness is not exclusive to quantum machine learning but is observed in overparametrized classical neural networks with large widths as well. The efficiency of large-width neural networks is substantiated by the neural tangent kernel theory, and similar principles apply to their quantum counterparts.
III Fault tolerant quantum computation (FTQC) algorithms

### v2

@@ 2.2 Quantum neural tangent kernel
$d\bar{}\hskip 1.00006ptz_{i;\delta}=d\bar{}\hskip 1.00006pt\varepsilon_{i;\delta}\equiv\frac{\partial z_{i;\delta}}{\partial\theta_{\ell}}\left(\mathcal{D}\theta_{\ell}\right)=\frac{dz_{i;\delta}}{d\theta_{\ell}}\left(-\eta\frac{\partial\mathcal{L}_{\mathcal{A}}}{\partial\theta_{\ell}}\right)=-\eta\sum_{j,\alpha,\ell}\varepsilon_{i^{\prime};\alpha}\frac{\partial\varepsilon_{i;\delta}}{\partial\theta_{\ell}}\frac{\partial\varepsilon_{i^{\prime};\alpha}}{\partial\theta_{\ell}}=-\eta\sum_{j,\alpha}\varepsilon_{i^{\prime};\alpha}K^{i,i^{\prime}}_{\delta,\alpha}\leavevmode\nobreak\ ,$ (5)
The variational angles $\theta_{\ell}$ are then updated step by step, labeled by $t$, according to the derivative of the loss function with respect to $\theta_{\ell}$. $\eta$ is defined as the learning rate, which is assumed to be a small (less than 1), positive number. The learning rate determines the size of the step that we take to update our parameters. One can calculate the update of the loss function $\mathcal{L}$ and prove that it decreases indefinitely every time we update our parameters, which is also a fact in classical machine learning [112]. In Eqn. (5) it is easy to show that the update of the output is mathematically in the exact same form as the update of the residual training error $d\bar{}\hskip 1.00006pt\varepsilon_{i;\delta}$. One then defines the quantum neural tangent kernel (QNTK):
$K^{i,i^{\prime}}_{\delta,\alpha}\equiv\sum_{\ell}\frac{\partial\varepsilon_{i;\delta}}{\partial\theta_{\ell}}\frac{\partial\varepsilon_{i^{\prime};\alpha}}{\partial\theta_{\ell}}\leavevmode\nobreak\ .$ (6)
@@
However, the final assumption is not always true. To wit, with the expression of the output $z_{i,\alpha},$ one can analytically calculate the QNTK and prove that the QNTK is highly fluctuating in the variational angles $\bm{\theta}$, rather than being a constant [87]. This fluctuation implies that we have obtained representation learning. In the terminology of the kernel method (NTK) of classical machine learning, it means that the feature map itself is changing. This indicates that we are extracting features from all of our training data to update our output, and we cannot obtain a closed-form solution for $\varepsilon(t)$ in this case.
@@
Figure 3: Demonstration of the evolution of the loss function in the continuous limit [87]. The right diagram depicts the closed-form solution when the QNTK remains constant. In contrast, the left diagram illustrates a scenario where the QNTK is nonlinear but nearly constant. In this case, representation learning is achieved, and the gradient descent equation can be addressed perturbatively.
An instance where the QNTK remains nearly constant is referred to as “lazy training”. This occurs when the variational angles are nearly unchanged [87, 33]: $\theta_{\ell}=\theta_{\ell}^{*}+\delta\varphi\leavevmode\nobreak\ .$ If $\theta_{\ell}$ remains approximately constant, indicating lazy training with minimal movement, the QNTK tends to be nearly constant as well. In such cases, the QNTK expression can be expanded with respect to $\theta_{\ell}$ for predictable dynamics, akin to classical scenarios [112].
@@
$\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu=\frac{L\operatorname{Tr}(\mathcal{O}_{i}^{2})}{N^{2}}\propto L\leavevmode\nobreak\ .$ (7)
@@
Nonetheless, this model remains useful, as even when not operating at the frozen limit, it provides an approximate means to gauge the speed of gradient descent. For example, to determine which variational ansatz runs the fastest without actually executing them, one can calculate their QNTK and select the one with the largest value. It also enables the estimation of operation time and the number of iterations required for the gradient descent program using Eqn. (7), facilitating hardware design improvements. Moreover, some late-time behaviors of QNTK also been recently captured in [138].
2.3 Quantum landscapes and barren plateaus
@@
The process of training parameters in the Quantum Machine Learning (QML) model often involves minimizing a loss function and traversing a non-convex landscape in search of its global minimum, as extensively discussed in [27]. Quantum landscape theory, as outlined in [9], seeks to comprehend the properties of QML landscapes and explores strategies for engineering them. Notably, local minima and barren plateaus are focal points of investigation within quantum landscape theory
@@
Akin to classical ML [27], the quantum loss landscape may contain numerous local minima. In a manner similar to classical cases, the overall non-convex optimization can become NP-hard [20]. Various approaches have been proposed to mitigate issues related to local minima. For instance, variable structure Quantum Neural Networks (QNNs), as introduced in [19, 81], dynamically adjust the model’s prior by expanding and contracting during optimization, transforming certain local minima into saddle points. Additionally, there is evidence of the overparametrization phenomenon in QML [73, 79]. In this scenario, a computational phase transition occurs as the optimization progresses, leading to the disappearance of spurious local minima when the number of parameters surpasses a critical value.
@@
Not only are local minima a concern in QML, but there is also the intriguing phenomenon of a barren plateau in quantum landscapes [27, 25, 8, 60]. A barren plateau is characterized by the loss landscape becoming, on average, exponentially flat concerning the problem size. In this scenario, the valley housing the global minimum exponentially diminishes with the problem size, forming what is known as a narrow gorge [9]. Consequently, exponential resources, such as an increased number of shots, are needed to traverse through the landscape. This impact on resource requirements complicates the QML algorithm’s complexity and can potentially undermine quantum speedup, as quantum algorithms typically aim to avoid the exponential complexity associated with classical algorithms.
@@
The barren plateau phenomenon was first studied in deep hardware-efficient QNNs [98], where they arise due to the high expressivity of the model [8, 27]. By making no assumptions about the underlying data, deep hardware-efficient architectures aim to solve a problem by being able to prepare a wide range of unitary evolutions. In other words, the prior over the hypothesis space is relatively uninformed. Barren plateaus in this unsharp prior are caused by ignorance or the lack of sufficient inductive bias, and therefore a means to avoid them is to input knowledge into the construction of the QNN - making the design of QNNs with good inductive biases for the problem at hand a key solution. Various strategies have been developed to address these barren plateaus, such as clever initialization [126], pretraining, and parameter correlation [60]. These are also examples of adding a sharper prior to one’s search over the over-expressive parameterizations of hardware-efficient QNNs. The authors of [28] studied the barren plateau problem for parameterized quantum circuits of the quantum tensor network architecture and show, for example, that classical optimization of these circuits can be exponentially more efficient than using a quantum computer.
@@
Other mechanisms have been linked to barren plateaus [27]. Simply defining a loss function based on a global observable (i.e., observables measuring all qubits) leads to barren plateaus even for shallow circuits with sharp priors [98], while local observables (those comparing quantum states at the single-qubit level) avoid this issue [98, 125]. The latter is not due to bad inductive biases but rather to the fact that comparing objects in exponentially large Hilbert spaces requires an exponential precision, as their overlap is usually exponentially small.
@@
While entanglement is one of the most important quantum resources for information processing tasks in quantum computers, it can also be detrimental for QML models [27]. QNNs (or embedding schemes) that generate too much entanglement also lead to barren plateaus [116, 95, 106]. Here, the issue arises when one entangles the visible qubits of the QNN (those that one measures at the QNN’s output) with a large number of qubits in the hidden layers. Due to entanglement, the information of the state is stored in non-local correlations across all qubits, and hence the reduced state of the visible qubits concentrates around the maximally mixed state. This type of barren plateau can be solved by taming the entanglement generated across the QNN.
2.3.1 Influence of background noise
@@
The existence of hardware noise during quantum computations is a distinctive feature of NISQ computing [27]. Despite this reality, many QML studies overlook noise in analytical calculations and numerical simulations, while still asserting near-term compatibility of the methods. It is imperative to consider the impact of hardware noise in QML analyses, especially for those aiming to achieve quantum advantage with currently available hardware [27].
@@
Noise introduces errors in the information as it propagates through a quantum circuit, particularly affecting deeper circuits with longer run-times [27]. This impact extends to all aspects of quantum models, including the dataset preparation scheme and circuits used for computing quantum kernels. In the case of Quantum Neural Networks (QNNs), noise can impede their trainability [128, 127], leading to noise-induced barren plateaus where relevant features are exponentially suppressed with increasing circuit depth. The consequence is a deformation of the model’s inductive bias and an effective reduction in the dimension of the quantum feature space. Despite the significant influence of quantum noise, its effects remain largely unexplored, especially regarding its impact on the classical simulability of QML models [39, 53, 27].
@@
Addressing issues induced by noise may necessitate [27]: (1) reducing hardware error rates, (2) implementing partial quantum error correction [24], or (3) utilizing QNNs that are relatively shallow (i.e., with depth growing sub-linearly in the problem size) [128], such as Quantum Convolutional Neural Networks (QCNNs). While error mitigation techniques [124, 38, 42] can enhance the performance of QML models in the presence of noise, they may not completely resolve noise-induced trainability issues [127]. Another approach involves engineering QML models with noise-resilient properties [80, 35, 115], such as ensuring that the position of minima remains unchanged despite noise.
@@
However, another study suggests that noise in VQAs might be beneficial for their performance. Drawing inspiration from classical machine learning, where noise is intentionally introduced to enhance the performance of the gradient descent algorithm, this phenomenon is akin to stochastic gradient descent. It has been observed that gradient descent tends to perform better under the influence of noise, as the noise helps in avoiding saddle points that could hinder the algorithm’s progress [69, 70]. In the absence of noise, the algorithm may converge directly to the saddle point. However, when noise is introduced, perturbations around the points prevent convergence towards these unstable fixed points.
@@
In the study [88], this discussion was extended to quantum noise. Through experiments, the researchers found that the noisy case outperformed the noiseless counterparts, providing a toy model illustration to support their argument. This suggests that, while implementing quantum error correction is crucial for making quantum devices fault-tolerant, in certain QML applications, eliminating all noise may not be necessary. As long as the noise remains below a certain threshold, it may prove not detrimental but rather helpful for QML programs.
2.3.2 Laziness
@@
$\theta_{\ell}(t+1)-\theta_{\ell}(t)=-\eta\frac{\partial\mathcal{L}}{\partial\theta_{\ell}}\equiv\delta\theta_{\ell}\leavevmode\nobreak\ .$ (9)
@@
The observation [98] is that if our variational ansatz is highly random, according to the $k$-design integral formula [111, 85, 84, 36], the derivative of the loss function is generally suppressed by the dimension of the Hilbert space $N$. This might lead to a situation where the variation of the loss function during gradient descent is very small, denoted as $\delta\mathcal{L}\equiv\mathcal{L}(t+1)-\mathcal{L}(t)\ll 1$ for the step $t$. An alternative theoretical term used to describe the quantum barren plateau, based on the large suppression in Eqn. (9), is laziness [86]. In the quantum context, the suppression comes from the dimension of the Hilbert space, while in the classical case, the suppression is associated with the width of classical neural networks.
@@
To be more precise, laziness refers to small $\delta\theta_{\mu}$, and the barren plateau refers to small $\delta\mathcal{L}$. The authors of [86] demonstrate that laziness may not necessarily imply the quantum barren plateau. This insight is derived from the perspectives of overparametrization theory and representation learning theory through the QNTK [87]. In the context of quantum neural networks, overparametrization refers to the condition $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu\approx\mathcal{O}(1)$ in Eqn.(7).
@@
The paper [86] emphasizes that in variational circuits with a sufficiently large number of trainable angles, gradient descent dynamics can still be efficiently executed, even in the presence of the exponential suppression of variational angle updates (laziness). On the other hand, the only limitation is the precision, where noises might affect the performance significantly. The authors also highlight that laziness is not exclusive to quantum machine learning but is observed in overparametrized classical neural networks with large widths as well. The efficiency of large-width neural networks is substantiated by the neural tangent kernel theory, and similar principles apply to their quantum counterparts.
3 Fault tolerant quantum computation (FTQC) algorithms